Determine the minimum volume of 0.50 mol/L calcium chloride solution required to precipitate 14 g of calcium carbonate when mixed with excess sodium carbonate


1
Expert's answer
2022-04-08T07:05:05-0400

CaCl2 + Na2CO3 = CaCO3 + 2NaCl

n(CaCO3) = m(CaCO3) / Mr(CaCO3) = 14 g / 100 g/mol = 0.14 mol

1 mol CaCl2 - 1 mol CaCO3

x mol CaCl2 - 0.14 mol CaCO3

x = 0.14 mol

V(sol) = n(CaCO3) / C(sol) = 0.14 mol / 0.50 mol/l = 0.28 l = 280 ml